OpenStudy (anonymous):

evaluate limit (lim x-> pi/2) pi-x/cos x

OpenStudy (helder_edwin):

just replace \(x=\pi/2\)

OpenStudy (helder_edwin):

got it?

OpenStudy (helder_edwin):

\[ \large \lim_{x\to\color{red}{\pi/2}}\frac{\pi-x}{\cos x}= \frac{\pi-\color{red}{\pi/2}}{\cos\color{red}{\pi/2}}= \]

OpenStudy (shubhamsrg):

am quite sure he means pi/2 -x in the numerator..please confirm..